A. How many grams of nickel metal will be deposited from a solution that contains Ni2+ions if a current of 1.21 A is applied for 48.0minutes.

B. How many grams of nickel metal will be deposited from a solution that contains Ni2+ions if a current of 0.664 A is applied for 38.0minutes. ______Grams

C. How many grams of cobalt metal will be deposited from a solution that contains Co2+ions if a current of 0.809 A is applied for 67.8minutes. ___________grams
Answer #1

a) Q = i*t

        = 1.21*48*60

       =3484.8 columbs

generally

2 *96500columbs current deposits -------------------------58.5 g of nickel

3484.8 columbs currrent deposits--------------------------------------?

the amount of nickel deposits = 3484.8*58.5/(2*96500)

                                              = 1.056g

b)

Q = i*t

        = 0.664*38*60

       =1513.92 columbs

generally

2 *96500columbs current deposits -------------------------58.5 g of nickel

1513.92 columbs currrent deposits--------------------------------------?

the amount of nickel deposits = 1513.92*58.5/(2*96500)

                                              = 0.458g

c)

Q = i*t

        = 0.809*67.8*60

       =3291.01columbs

generally

2 *96500columbs current deposits -------------------------58.9g of cobalt

3291.01columbs currrent deposits--------------------------------------?

the amount of cobalt deposits = 3291.01*58.9/(2*96500)

                                              = 1.004g